random231 (random231):

if y is a function of x. and log(x+y)-2xy=0 what is the value of f'(0)?

OpenStudy (kc_kennylau):

\[\begin{array}{rcl} \ln(x+y)-2xy&=&0\\ \ln(x+y)&=&2xy\\ \frac{1+\frac{dy}{dx}}{x+y}&=&2y+2\frac{dy}{dx}\\ \end{array}\]

OpenStudy (kc_kennylau):

then i have no idea xP

random231 (random231):

i did upto this.

OpenStudy (perl):

theres a mistake in your derivative

OpenStudy (perl):

log(x+y)-2xy=0 log (x+y) = 2xy (1 + y' ) / (x+y) = 2*1*y + 2*x*y'

OpenStudy (kc_kennylau):

what was the mistake?

OpenStudy (perl):

your right side, should be 2y + 2xy'

OpenStudy (kc_kennylau):

oh i see

OpenStudy (perl):

now substitute y'(0) , , given that x = 0, and y = (you can solve for this)

OpenStudy (perl):

mustafa made the same mistake -sigh-

OpenStudy (anonymous):

at x=0 y=1

OpenStudy (perl):

log(x+y)-2xy=0 log (x+y) = 2xy (1 + y' ) / (x+y) = 2*1*y + 2*x*y'

random231 (random231):

okay

OpenStudy (perl):

when x = 0, then log( 0 + y ) - 2(0)*y = 0 log y - 0 = 0 log y = 0 y = 1. now plug in , x = 0, y = 0, and solve for y'(0)

OpenStudy (perl):

(1 + y'(0) ) / (0+1) = 2*1*1 + 2*0*y'(0)

OpenStudy (perl):

( 1 + y'(0) ) /1 = 2 y'(0) = 1

random231 (random231):

thanks